If 3^x=17, then x equals what?
Answer by RAY100(1637)

when variable is an exponent, normally use log to solve.
.
Remember loga^b = b*loga
.
3^x = 17
log(3^x) = log(17)
x*log (3) =log(17)
x=log(17)/log(3)
x=2.5789
.
check,,,3^2.5789 =17,,,ok